OpenStudy (anonymous):

How do I do this? If x + 2y = 6 and 3x + y = 4, then 4x + 3y =

OpenStudy (phi):

add your 2 equations

OpenStudy (phi):

you could solve for x and then y (or vice versa), but notice what happens if you just add the 2 equations together : x + 2y = 6 3x + y = 4 ---------- 4x + 3y = 10